### v2

Abstract In this work, we introduce a partitioned optimization framework (POf) to ease the solving process for optimization problems for which fixing some variables to a tunable value removes some difficulties. The variables space may be continuously partitioned into subsets where these variables are fixed to a given value, so that minimizing the objective function restricted to any of the partition sets is easier than minimizing the objective function over the whole variables space. Then our methodology consists in searching for the partition set minimizing the minimum of the objective function restricted to this set. Indeed, a local solution to the original problem is given by computing this partition set and selecting the minimizer of the objective function restricted to this set. We first formalize this framework and study its theoretical guarantees, and we discuss about a derivative-free optimization (DFO) algorithm able to seek for an optimal partition set. Then we focus on a class of composite optimization problems that may be solved using the POf, and we illustrate in details our methodology on four problems from this class. Finally, we highlight the gain in numerical performance provided by the POf on large-dimensional composite problems by comparing it to two popular DFO solvers.
Data availability statement: This manuscript has no associated data.
Remark: This document consists in our work displayed in a dedicated template designed for the reviewing process. Another version written in a generic template closer to an actual paper is available at https://arxiv.org/abs/2407.05046. The content is exactly the same on both versions; only the displaying differs.
{GDacknowledgements}
This work was supported by the first NSERC discovery grant RGPIN-2020-04448 of Audet.
[…]
1 Introduction
In this work, we consider the optimization problem
$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{ini}}$)
where $\mathbb{Y}$ is a (possibly infinite-dimensional) normed space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}$ is a (possibly discontinuous) function with nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$, and $\Omega\subseteq\mathbb{Y}$ intersects $Y$. We introduce a partitioned optimization framework (POf) to solve Problem ($\mathbf{P}_{\texttt{ini}}$) in cases where $\mathbb{Y}$ may be partitioned in a way that the restriction of the problem to any partition set is significantly simpler to solve than the problem itself. Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 summarizes our contribution and outlines this work.
Notation: We denote by $\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$$\overline{\mathbb{N}}\triangleq\mathbb{N}\cup\{+\infty\}$$\mathbb{R}^{*}\triangleq\mathbb{R}\setminus\{0\}$ and ¯¯$\overline{\mathbb{N}}^{*}\triangleq\overline{\mathbb{N}}\setminus\{0\}$. Consider two normed spaces $\mathbb{A}$ and $\mathbb{B}$. A set $A\subseteq\mathbb{A}$ is said to be ample [2] if $A\subseteq\mathrm{cl}(\mathrm{int}(A))$ (where $\mathrm{cl}$ and $\mathrm{int}$ denote respectively the closure and interior operators), and to be a continuity set of a function $f:\mathbb{A}\to\overline{\mathbb{R}}$ if $f(a)$ if finite for all $a\in A$ and $f_{|A}$ is continuous. For all collections $(A_{i})_{i=1}^{N}$ of subsets of $\mathbb{A}$$N\in\overline{\mathbb{N}}^{*}$, their union is denoted by $\sqcup_{i=1}^{N}A_{i}$ when the sets are pairwise disjoint. For all sequences $(b^{k})_{k\in\mathbb{N}}\in\mathbb{B}^{\mathbb{N}}$, we denote by $\mathrm{acc}((b^{k})_{k\in\mathbb{N}})$ the set of all accumulation points of $(b^{k})_{k\in\mathbb{N}}$. For all functions $g:\mathbb{A}\to\mathbb{B}$ and all points $a\in\mathbb{A}$, we denote by $\mathcal{ACC}(g;a)$ the union of $\mathrm{acc}((g(a^{k}))_{k\in\mathbb{N}})$ over all sequences $(a^{k})_{k\in\mathbb{N}}\in\mathbb{A}^{\mathbb{N}}$ converging to $a$. For all functions $f:\mathbb{A}\to\overline{\mathbb{R}}$ and all sets $A\subseteq\mathbb{A}$, we denote by $\overline{f_{A}}:\mathbb{A}\to\overline{\mathbb{R}}$ the extreme barrier function [3] mapping all $a\in\mathbb{A}$ to $f(a)$ if $a\in A$ and to +$\infty$+ ∞ if $a\notin A$. In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{ini}}$).
1.1 Illustrative example with a composite greybox problem
This section illustrates the POf with no constraints ($\Omega\triangleq\mathbb{Y}$), two variables ($\mathbb{Y}\triangleq\mathbb{R}^{2}$), and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ (y_{1},y_{2})&\mapsto&(y_{2}-\sigma(y_{1}))^{2}+\varepsilon(y_{1})\end{array}\right.$
shown in Figure 1 (this example is formally solved in Section 4.1). Note that $\varphi$ is explicitly defined with respect to the functions $\sigma$ and $\varepsilon$, while $\sigma$ and $\varepsilon$ are blackboxes since their expressions are not provided. A possible application of the POf is to exploit partial information in similar contexts.
Figure 1: The function $\varphi$ from Section 1.2, minimized efficiently via the POf. For readability, at each $y_{1}\in[-4,4]$ the two graphs actually plot only the values of the quadratic $y_{2}\mapsto\varphi(y_{1},y_{2})$ for $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$.
Minimizing $\varphi$ over $\mathbb{Y}$ is difficult, remark that $\varphi$ has a simple expression when $y_{1}$ is fixed: for all $y_{1}\in\mathbb{R}$, the function $y_{2}\mapsto\varphi(y_{1},y_{2})$ is quadratic. The first step of the POf exploits this remark by partitioning $\mathbb{Y}$ as $\mathbb{Y}\triangleq\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$, where $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, so that minimizing $\varphi$ over $\mathbb{Y}(x)$ would be simple: in this case, we directly observe that the solution is $(x,\sigma(x))$, with objective function value $\varepsilon(x)$. Then, the POf defines an oracle function $\gamma:\mathbb{X}\to\mathbb{Y}$ that, given any partition index $x\in\mathbb{X}$, computes the minimizer of $\varphi$ over $\mathbb{Y}(x)$; in this case $\gamma(x)\triangleq(x,\sigma(x))$ is analytically tractable for all $x\in\mathbb{X}$. The second step of the POf calculates an optimal partition set index $x\in\mathbb{X}$ minimizing the minimizer of $\varphi$ over the set $\mathbb{Y}(x)$, that is, minimizing $\Phi(x)\triangleq\varphi(\gamma(x))$ (and, in this case, $\Phi(x)=\varepsilon(x)$). This problem is solved by an usual blackbox optimization approach, where we compute $\gamma(x)$ on-the-fly during the optimization process at points $x\in\mathbb{X}$ we seek to evaluate. Finally, the last step of the POf recovers a solution to Problem ($\mathbf{P}_{\texttt{ini}}$), given by $\gamma(x^{*})$, from an optimal partition set $x^{*}$ (in this example, $x^{*}\triangleq 0$ and $\gamma(x^{*})\triangleq(0,0)$). For completeness, define also the index function $\chi$, defined for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$ as the unique index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. In this context, $\chi(y)\triangleq y_{1}$ for all $y\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
1.2 Motivation for this work
This work originates from the corresponding author’s PhD thesis [9, Chapters 5 and 6] (in French). The initial motivation was to solve an optimal control problem [9, Chapter 7] with discontinuities, which justifies $\mathbb{Y}$ being possibly infinite-dimensional and $\varphi$ being possibly discontinuous. An early version of [9, Chapter 7], without the theory supporting the POf, appears in a technical report [10]. In all cases considered in [10]$\mathbb{Y}$ is an infinite-dimensional functional space and $\varphi$ is discontinuous, while the POf identifies a partition of $\mathbb{Y}$ such that $\dim(\mathbb{X})\leq 4$ and the problem to minimize $\varphi_{|\mathbb{Y}(x)}$ for any $x\in\mathbb{X}$ is an usual optimal control problem that may be solved by most dedicated techniques.
The POf allows to solve several problems that share similarities with the example from Section 1.1. What matters is that a well-chosen partition of $\mathbb{Y}$ allows for a significant simplification of the problem, so that a global solution to the problem restricted to any partition set is tractable (analytically of through a dedicated optimization algorithm) whenever one needs to calculate it. Then the problem is solved by seeking for the optimal partition set index, which is done as an usual blackbox problem (likely low-dimensional even when $\mathbb{Y}$ is infinite-dimensional). A possible field list of applications reads as composite greybox problem where the inputs of the blackbox components of the problem depend only on a few nonlinear combinations of all variables, as discussed in details in Section 4. Others possible applications are completely explicit problems (admitting no blackbox components) that consist in two distinct sub-problems that are connected through a few variables, or are discontinuous but the discontinuities are affected by a few variables only, or become smooth when a constraint linking the variables is added. In particular, our initial motivational problem [10] matches all these examples.
1.3 Contribution and outline of this work
Our contribution is threefold. First, we formalize the POf and we prove that an optimal partition set index allows to recover a solution to Problem ($\mathbf{P}_{\texttt{ini}}$) under a minimal assumption. Second, we provide an algorithm that computes an optimal partition set under an additional assumption. Third, we empirically measure the performance gain provided by the POf through numerical experiment, by comparing this approach to two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly.
Let us stress that ensuring the accessibility of the oracle function $\gamma$ is not within our scope. For ease of presentation and theoretical analysis, we design the POf as if we possess an idealized optimization method that may, for all partition set indices $x$, compute a global minimizer of the restriction $\varphi_{|\mathbb{Y}(x)\cap\Omega}$. Then $\gamma$ consists in the function that, given any $x$, calls this method and returns its output. Section 6 discusses about this restrictive setting and how we plan to significantly relax it in a future work. Similarly, how to find a relevant partition of $\mathbb{Y}$ in general is not within our scope either, since this is problem-dependent. Nevertheless, Section 4 shows how to find such a partition in a class of composite problems, and [9, Chapter 7] and [10] provide a similar guideline in a class of optimal control problems.
This work is organized as follows. A literature review is provided in Section 2. Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it formalizes our strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) using the POf in Theorem 2. Our third goal is considered in Sections 4 and 5. Section 4 highlights how Theorem 2 allows to solve some composite optimization problems, and shows that numerical observations match our theoretical claims. Section 5 alters the problems from Section 4 into more challenging versions with a large dimension, and shows that solving these problems with two DFO solvers from the literature (solving the non-reformulated problems) is less efficient than our strategy. Section 6 discusses about the POf and its possible extensions.
2 Literature review
The POf is not suited for blackbox optimization problems but is related to greybox optimization [1]. Indeed, choosing a relevant partition for $\mathbb{Y}$ requires information on the problem, and for all $x\in\mathbb{X}$, ensuring the tractability of $\gamma(x)$ requires that the restriction $\varphi_{|\mathbb{Y}(x)\cap\Omega}$ is simple enough to minimize. Nevertheless, most DFO algorithms [22] tackling greybox problems are initially designed for blackbox contexts, because fine-tuning strategies to greybox cases are problem-dependent. Then, we compute an optimal partition set using an algorithm with a covering step [2] since to our best knowledge only these algorithms are guaranteed to return a local solution in discontinuous contexts (under a mild assumption). We provide ideas in Remark 2 to specialize an algorithm to the context of the POf.
Solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly with an algorithm from derivative-free optimization (DFO) is likely inefficient, since the dimension of $\mathbb{Y}$ may be high while DFO algorithms usually handle up to a few dozens of variables [5, Section 1.4]. For example, both the NOMAD [6] and PRIMA [30] solvers state a critical value of $50$ variables. Some dimension-selection techniques [17] handle large-scale problems with active subspaces [14, 29], or via random directions [12, 19, 26]. These techniques likely outperform the POf when the objective function has a small sensitivity with respect to most of the variables, but likely fail on cases where the objective function has a small sensitivity with respect to each variables individually but has a high sensitivity with respect to all variables jointly. For example, this happens in some optimal control problems [9, Chapter 7]: an alteration of any single state variable has a small impact, but an alteration of all state variables together significantly alters the objective value.
The POf is related to parametric optimization [7, 28], which studies problems where the objective function and the feasible set depend on a tunable parameter. In our case, the parametric problem would be the minimization of $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem would be the partition set index $x\in\mathbb{X}$. Any problem fitting in any of the two frameworks may fit in the other. The local continuity of $\gamma$ in a discontinuous context is studied in [18], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours. Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{ini}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem (the index of the associated partition set of $\mathbb{Y}$, which results from the chosen partition of $\mathbb{Y}$) is left to the user.
The POf also shares similarities with bilevel optimization [15], which describes problems in which some of the variables are fixed as a solution to a nested subproblem. Any problem expressed in any of the frameworks may be transferred to the other. However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{ini}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure. Some references study bilevel nonsmooth problems [25] and apply DFO algorithms [16] but, to the best of our knowledge, none provides results similar to ours.
The terminology partitioned optimization also describes the DFO methods [20, 24]. However, these methods consider a discrete partition which do not simplifies the restricted problem, as they partition an hyperrectangle into smaller ones and recursively partition the most promising. Then, we believe that our continuous and structure-simplifying partition fits in the same terminology without confusion.
The POf may also be related to optimization on manifold [11], which considers problems with a constraint inducing an explicit manifold on the variables space. This happens in our case to the problem restricted to any partition set, when each partition set of $\mathbb{Y}$ is a manifold. In practice, when $\mathbb{Y}$ has a finite dimension, it is likely that $\mathbb{Y}$ is indeed partitioned into a continuum of manifolds.
3 Formal framework and related theoretical results
This section considers Problem ($\mathbf{P}_{\texttt{ini}}$) as stated in the beginning of the paper. That is, the objective function $\varphi$ is not limited to its expression from the introductory motivating example. The POf is formalized in Section 3.1, its properties follow in Section 3.2, and their proofs are given in Section 3.3.
3.1 Formal definition of the POf and of the algorithm we consider to exploit it
First, let us formalize the POf in Framework 1.
Framework 1 (partitioned optimization framework (POf)).
Let $\mathbb{X}$ be a finite dimensional normed space, and $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ be a partition such that the subproblem associated to Problem ($\mathbf{P}_{\texttt{ini}}$) at $x$,
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{sub}}(x)$)
admits a nonempty set of global solutions denoted by $\Gamma(x)$ for all $x\in X\triangleq\{x\in\mathbb{X}:\mathbb{Y}(x)\cap Y\cap\Omega\neq\emptyset\}$. Define an oracle function $\gamma:X\to Y$, satisfying $\gamma(x)\in\Gamma(x)$ for all $x\in X$ and assumed tractable. Define also the index function $\chi:\mathbb{Y}\to\mathbb{X}$ that maps all $y\in\mathbb{Y}$ to the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. Finally, define the reformulation of Problem ($\mathbf{P}_{\texttt{ini}}$),
$\underset{\begin{array}[]{c}x\in\mathbb{X}\end{array}}{\operatorname*{minimize}}\quad\Phi(x),\qquad\mbox{where}\qquad\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\overline{\mathbb{R}}\\ x&\mapsto&\left\{\begin{array}[]{ll}\varphi(\gamma(x))&\mbox{if}\leavevmode\nobreak\ x\in X,\\ +\infty&\mbox{otherwise}.\end{array}\right.\end{array}\right.$ ($\mathbf{P}_{\texttt{ref}}$)
Second, we introduce Algorithm 1 to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$). We consider the cDSM: the usual direct search method (DSM[3, 5, 13] from DFO, enhanced with a covering step [2]. The covering step ensures that Algorithm 1 returns local solutions, even in the presence of discontinuities. The cDSM is the first scheme discussed in [2], but this papers also fits a covering step into any algorithm. Then we could use any other DFO method instead of the DSM. Only the covering step matters in our analysis.
Algorithm 1 cDSM [2, Algorithm 1] solving Reformulation ($\mathbf{P}_{\texttt{ref}}$).
Initialization:
select $(x^{0},\delta^{0})\in X\times\mathbb{R}_{+}^{*}$ the initial (incumbent solution, poll radius) couple; select $(\lambda,\upsilon)\in{]}0,1{[}\times[1,+\infty{[}$ the poll radius shrinking and expanding parameters; set $\mathcal{H}^{0}\triangleq\emptyset$ the initial trial points history;
for $k\in\mathbb{N}$ do:
     covering step:
      set $\mathcal{D}_{\tt{C}}^{k}\triangleq\operatorname*{argmax}_{\lVert{d}\rVert\leq 1}\mathrm{dist}(x^{k}+d,\mathcal{H}^{k})$; set $\mathcal{T}_{\tt{C}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{C}}^{k}$ and $t_{\tt{C}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{C}}^{k})$; if $\Phi(t_{\tt{C}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{C}}^{k}$ and $\mathcal{T}_{\tt{S}}^{k}=\mathcal{T}_{\tt{P}}^{k}\triangleq\emptyset$ and go to the update step;
     search step:
      select $\mathcal{D}_{\tt{S}}^{k}\subseteq\mathbb{X}$ empty or finite; if $\mathcal{T}_{\tt{S}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{S}}^{k}$ is nonempty, then select $t_{\tt{S}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{S}}^{k})$; if moreover $\Phi(t_{\tt{S}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{S}}^{k}$ and $\mathcal{T}_{\tt{P}}^{k}\triangleq\emptyset$ and go to the update step;
     poll step:
      select $\mathcal{D}_{\tt{P}}^{k}\subseteq\mathbb{X}$ a positive basis of length $\delta^{k}$; set $\mathcal{T}_{\tt{P}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{P}}^{k}$; select $t_{\tt{P}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{P}}^{k})$; if $\Phi(t_{\tt{P}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{P}}^{k}$, otherwise set $t^{k}\triangleq x^{k}$;
     update step:
      set $(x^{k+1},\delta^{k+1})\triangleq(t^{k},\upsilon\delta^{k})$ if $t^{k}\neq x^{k}$ or $(x^{k+1},\delta^{k+1})\triangleq(x^{k},\lambda\delta^{k})$ if $t^{k}=x^{k}$; set $\mathcal{H}^{k+1}\triangleq\mathcal{H}^{k}\cup\mathcal{T}_{\tt{C}}^{k}\cup\mathcal{T}_{\tt{S}}^{k}\cup\mathcal{T}_{\tt{P}}^{k}$.
Remark 1.
Sections 4 and 5 instantiate Algorithm 1 as follows. The point $x^{0}$ depends on the instance, while we set $\delta^{0}\triangleq 1$. We choose $(\lambda,\upsilon)\triangleq(\frac{1}{2},1)$ in Sections 4.14.24.35.1 and 5.2, and $(\lambda,\upsilon)\triangleq(\frac{1}{2},2)$ in Section 5.3, and $(\lambda,\upsilon)\triangleq(\frac{3}{4},2)$ in Sections 4.4 and 5.4. The search step is skipped ($\mathcal{D}_{\tt{S}}^{k}\triangleq\emptyset$) and the poll step selects random orthogonal positive bases. The algorithm iterates until $\delta^{k}<1E^{-10}$.
Remark 2.
A sensitivity analysis of Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with respect to $x$ via the generalized implicit function theorem [27] locally describes the oracle function $\gamma$. This may provide a descent direction for $\Phi$ emanating from $x$. Also, a surrogate [8] of the problem exists when $\gamma(x)$ is obtained numerically for all $x\in X$, by approximating $\gamma(x)$ as the output of a method solving ($\mathbf{P}_{\texttt{sub}}(x)$) on low precision.
3.2 Theoretical properties of the POf and Algorithm 1
To study the theoretical properties of the POf, we define first the notion of generalized local solutions, hinted in [2] but not formally stated, that is suited for minimization problems with a possibly nonclosed feasible set and a possibly not lower semicontinuous objective function. We introduce it in Definition 1, for a generic problem since we will apply it to either Problem ($\mathbf{P}_{\texttt{ini}}$) and Reformulation ($\mathbf{P}_{\texttt{ref}}$).
Definition 1 (Generalized local solutions).
Consider a normed space $\mathbb{A}$, and $A\subseteq\mathbb{A}$ and $f:\mathbb{A}\to\overline{\mathbb{R}}$. Any point $x^{*}\in\mathrm{cl}(A)$ is said to be a generalized local solution to the problem
$\underset{\begin{array}[]{c}a\in\mathbb{A}\end{array}}{\operatorname*{minimize}}\quad f(a)\quad\mbox{subject to}\quad a\in A$
if there exists a neighbourhood $\mathcal{A}^{*}\subseteq\mathbb{A}$ of $a^{*}$ such that $\inf\overline{f_{A}}(\mathcal{A}^{*})=\inf\mathcal{ACC}(\overline{f_{A}},a^{*})$.
Assumption 1.
[…]
Theorem 1.
In Framework 1 under Assumption 1, the next three statements hold:
  1. a)
    for all global solutions $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
  2. b)
    for all local solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
  3. c)
    for all generalized local solutions $x^{*}\in\mathrm{cl}(X)$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
Our second result formalizes a strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) under an additional assumption.
Assumption 2.
[…]
Theorem 2.
In Framework 1 under Assumptions 1 and 2, Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$, the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$). Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, if $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
We prove Theorems 1 and 2 in Section 3.3, and we discuss Assumptions 1 and 2 in Section 6.1. To conclude this section, remark that Theorem 1 is an inner property of the POf, while Theorem 2 consists in an algorithmic strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) via the POf. Similarly, Assumption 1 is a minimal assumption on the POf, while Assumption 2 is an additional requirement required only to ensure that Algorithm 1 works as expected. Remark also that any DFO algorithm could be used instead of Algorithm 1, provided that it possesses a convergence analysis similar to those of the cDSM (stated in Proposition 1 below).
3.3 Proofs of Theorems 1 and 2
First, we state some lemmas related to the POf and Assumption 1 in Section 3.3.1. Second, we prove Theorem 1 in Section 3.3.2. Third, in Section 3.3.3 we settle the ground for the proof of Theorem 2: this section recalls the convergence analysis of Algorithm 1 from [2, Theorem 1] as Proposition 1 and proves that it is applicable when Assumption 2 holds. Finally, we prove Theorem 2 in Section 3.3.4.
3.3.1 technical lemmas on the problem, the framework and its minimal assumption
This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{ini}}$), the POf, and Assumption 1. Lemma 1 is related to our notation $\overline{\varphi_{\Omega}}$ and $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and Lemma 2 states a basic property of the POf. Lemma 3 allows us, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{ini}}$). Lemmas 4 and 5 focus on the set $\mathcal{ACC}(\gamma;x)$ involved in Theorems 1.c) and 2. In particular, we prove Theorem 1.c) by using Lemma 5 to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
Lemma 1.
For all $y\in\mathbb{Y}$ and all $\mathcal{Y}\subseteq\mathbb{Y}$ neighbourhood of $y$, it holds that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$.
Proof.
Let $y\in\mathbb{Y}$ and let $\mathcal{Y}\subseteq\mathbb{Y}$ be a neighbourhood of $y$. Let $z\in\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and $(y^{k})_{k\in\mathbb{N}}\in\mathcal{Y}^{\mathbb{N}}$ be a sequence converging to $y$ and such that $(\overline{\varphi_{\Omega}}(y^{k}))_{k\in\mathbb{N}}$ converges to $z$. Then, for all $k\in\mathbb{N}$ it holds that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq\overline{\varphi_{\Omega}}(y^{k})$, so we deduce that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq z$. ∎
Lemma 2.
[…]
Proof.
For all pairs $(x_{1},x_{2})\in X^{2}$ with $x_{1}\neq x_{2}$, it holds that $\mathbb{Y}(x_{1})\cap\mathbb{Y}(x_{2})=\emptyset$, and $\gamma(x_{1})\in\mathbb{Y}(x_{1})$ and $\gamma(x_{2})\in\mathbb{Y}(x_{2})$. It follows that $\gamma(x_{1})\neq\gamma(x_{2})$. Hence $\gamma$ is an injection from $X$ to $\gamma(X)$, and thus a bijection. Also, for all $x\in X$, the inclusion $\gamma(x)\in\mathbb{Y}(x)$ raises $\chi(\gamma(x))=x$, and for all $y\in\gamma(X)$, there exists a unique $x\in X$ satisfying $y=\gamma(x)\in\mathbb{Y}(x)$, so $\chi(y)=x$ and then $\gamma(\chi(y))=\gamma(x)=y$. ∎
Lemma 3.
Under Assumption 1, for all $x\in\mathbb{X}$ and all $\mathcal{X}\subseteq\mathbb{X}$ neighbourhood of $x$, the set $\mathbb{Y}(\mathcal{X})$ is a neighbourhood of all $y\in\mathbb{Y}(x)$.
Proof.
Consider that Assumption 1 holds. Let $x\in\mathbb{X}$, and $\mathcal{X}\subseteq\mathbb{X}$ be a neighbourhood of $x$, and $y\in\mathbb{Y}(x)$. Then $\chi(y)=x\in\mathcal{X}$, and by continuity of $\chi$ there exists an open set $\mathcal{Y}\subseteq\mathbb{Y}$ containing $y$ and such that $\chi(\mathcal{Y})\subseteq\mathcal{X}$, so $y\in\mathcal{Y}\subseteq\mathbb{Y}(\mathcal{X})$. ∎
Lemma 4.
Under Assumption 1, for all $x\in\mathrm{cl}(X)$ and all $(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converging to $x$, it holds that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})\subseteq\mathcal{ACC}(\gamma;x)\subseteq\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$ and $\mathcal{ACC}(\gamma;x)$ is finite.
Proof.
Consider that Assumption 1 holds. Let $x\in\mathrm{cl}(X)$ and let $\mathcal{X}\subseteq\mathbb{X}$ be the neighbourhood of $x$ given by Assumption 1, so $\mathcal{I}\triangleq\{i\in\llbracket 1,N\rrbracket:X_{i}\cap\mathcal{X}\neq\emptyset\}$ is finite. For all $i\in\mathcal{I}$, by uniform continuity of $\gamma_{|X_{i}}$, the set $\cup_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X_{i}^{\mathbb{N}}}\leavevmode\nobreak\ \mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})$ is a singleton, and its unique element denoted by $y_{i}$ satisfies $y_{i}\in\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$ since $\gamma(X)\subseteq Y\cap\Omega$ and by continuity of $\chi$. Then, $\mathcal{ACC}(\gamma;x)=\{y_{i},\leavevmode\nobreak\ i\in\mathcal{I}\}$ is a finite subset of $\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$. Now, let $(x^{k})_{k\in\mathbb{N}}\in(X\cap\mathcal{X})^{\mathbb{N}}$ be a sequence converging to $x$ and, for all $i\in\llbracket 1,N\rrbracket$, define $K_{i}\triangleq\{k\in\mathbb{N}:x^{k}\in X_{i}\}$. Then, +$\mathcal{I}_{+}\triangleq\{i\in\mathcal{I}:K_{i}$ is infinite}}\}} is nonempty since we have $\mathbb{N}=\sqcup_{i\in\llbracket 1,N\rrbracket}K_{i}=\sqcup_{i\in\mathcal{I}}K_{i}$. We deduce that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})=\{y_{i},\leavevmode\nobreak\ i\in\mathcal{I}_{+}\}\subseteq\mathcal{ACC}(\gamma;x)$. ∎
Lemma 5.
[…]
Proof.
Consider that Assumption 1 holds and let $x\in\mathrm{cl}(X)$. By Lemma 4, it holds that
[…]
so it follows that $\inf\mathcal{ACC}(\Phi;x)\geq\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ as claimed. ∎
3.3.2 Proof of Theorem 1
Consider that Framework 1 is applicable and that Assumption 1 holds.
Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists). For all $y\in\Omega$, it holds that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. This proves Theorem 1.a).
Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(\mathbb{Y}(\mathcal{X}^{*}))\subseteq\mathcal{X}^{*}$. Thus, for all $y\in\mathbb{Y}(\mathcal{X}^{*})\cap\Omega$, it follows that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. This proves Theorem 1.b).
Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so $\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ is nonempty. Select an element $y^{*}$ from it. First, there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\inf\mathcal{ACC}(\Phi;x^{*})=\inf\Phi(\mathcal{X}^{*})$, and $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $y^{*}$ by Lemma 3. Second, $y^{*}$ satisfies $\inf\mathcal{ACC}(\Phi;x^{*})\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ by Lemma 5. Third, $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\Phi(\mathcal{X}^{*})$ holds since all $y\in\mathbb{Y}(\mathcal{X}^{*})$ satisfy $\overline{\varphi_{\Omega}}(y)\geq\Phi(\chi(y))\geq\inf\Phi(\mathcal{X}^{*})$. We deduce that $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$, and then $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))=\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ via Lemma 1. Thus $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.c). ∎
3.3.3 Convergence analysis of Algorithm 1 and proof of validity under Assumption 2
This section starts with a rewording of the convergence analysis of Algorithm 1 from [2, Theorem 1], as Proposition 1. Then, Propositions 2 and 3 show that all its assumptions are all met in our context.
Proposition 1 (convergence analysis of Algorithm 1).
Assume that $\Phi$ is bounded below with bounded sublevel sets and that $X$ admits a partition into $N\in\overline{\mathbb{N}}^{*}$ ample continuity sets of $\Phi$. Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, if $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
Proposition 2.
if Assumptions 1 and 2 hold, then $\Phi$ is bounded below and has its sublevel sets bounded, and $X_{i}$ is an ample continuity set of $\Phi$ for each $i\in\llbracket 1,N\rrbracket$.
Proof.
First, $\inf\Phi(\mathbb{X})=\inf\Phi(X)=\inf\varphi(\gamma(X))>{-}\infty$ holds by assumption, so $\Phi$ is bounded below. Second, let $\alpha\in\mathbb{R}$. For all $x\in\mathbb{X}$, the inclusion $x\in\Phi^{-1}({]}{-}\infty,\alpha])$ is equivalent to $\Phi(x)\leq\alpha$, and then to $\varphi(\gamma(x))\leq\alpha$, and then to $\gamma(x)\in\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$, and then to $x\in\chi(\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha]))$ by Lemma 2. Thus $\Phi^{-1}({]}{-}\infty,\alpha])$ is bounded, since by assumption $\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$ is bounded and $\chi(\mathcal{Y})$ is bounded for all $\mathcal{Y}\subseteq\mathbb{Y}$ bounded. Last, consider the partition of $X$ given by Assumption 2. For all $i\in\llbracket 1,N\rrbracket$$X_{i}$ is ample and $\Phi_{|X_{i}}=(\varphi_{|\gamma(X_{i})})\circ(\gamma_{|X_{i}})$ is continuous as a composition of continuous functions. ∎
Proposition 3.
Under Assumption 1, for all $x\in X$, if $\mathcal{ACC}(\gamma;x)\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x)$, then $\Phi$ is lower semicontinuous at $x$.
Proof.
Consider that Assumption 1 holds, let $x\in X$ and assume that $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x)\subseteq\Omega$. Lemma 5 claims that $\inf\mathcal{ACC}(\Phi;x)\geq\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$. Moreover, for all $y\in\mathcal{ACC}(\gamma;x)$, it holds that $\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)\geq\inf\mathcal{ACC}(\varphi;y)\geq\varphi(y)$ by lower semicontinuity of $\varphi$ at $y$, and that $\varphi(y)\geq\varphi(\gamma(\chi(y)))=\Phi(x)$ since $y\in\Omega$ and $\chi(y)=x\in X$ so $\gamma(\chi(y))$ exists. Then, we deduce that $\inf\mathcal{ACC}(\Phi;x)\geq\Phi(x)$, which proves the claim. ∎
3.3.4 Proof of Theorem 2
Consider that Framework 1 is applicable and that Assumptions 1 and 2 hold. Let $(x^{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$). Proposition 2 shows that Proposition 1 is applicable, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$), and then $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.c). Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, Proposition 3 claims that $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$, then $\Phi$ is lower semicontinuous at $x^{*}$. In this last case, $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.b). This completes the proof of Theorem 2. ∎
4 Application of the POf to composite optimization problems
In this section, we focus on a class of composite optimization problems (see [21, 23] and many references therein, or [22, Section 5]) where $\Omega$ is described by smooth constraints and $\varphi$ is defined by
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\overline{\mathbb{R}}\\ y&\mapsto&\widetilde{\varphi}(y,\sigma(f(y)))+\varepsilon(f(y)),\end{array}\right.$
where $\widetilde{\varphi}:\mathbb{Y}\times\mathbb{R}\to\mathbb{R}$ is smooth and easy to minimize and has an explicit definition, and $f:\mathbb{Y}\to\mathbb{X}$ is smooth with explicit definition (and $\dim(\mathbb{X})\in\mathbb{N}^{*}$ small), and $\sigma:\mathbb{X}\to\mathbb{R}$ and $\varepsilon:\mathbb{X}\to\overline{\mathbb{R}}$ are blackboxes (that is, their expression is unknown). We say that $\varphi$ is a greybox, since its explicit expression depends on some blackbox components. The POf (via the strategy formalized in Theorem 2) allows to solve such composite greybox problems, as shown below. This section solves four examples using this strategy.
In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{ini}}$). Indeed, $f(y)=x$ is a smooth equality constraint and $\sigma(f(y))=\sigma(x)$ and $\varepsilon(f(y))=\varepsilon(x)$ are constant. Accordingly, we partition $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$. Then, either Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed. Thus, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise. Solving it is manageable with the cDSM and, according to Theorem 2, this allows to recover solutions to the large-dimensional Problem ($\mathbf{P}_{\texttt{ini}}$). Remark also that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds.
[…]
4.1 Mono-variable noise
Consider again the example from Section 1.1, defined via $\mathbb{Y}=Y\triangleq\mathbb{R}^{2}$, and $\Omega\triangleq\mathbb{Y}$, and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ (y_{1},y_{2})&\mapsto&\left(y_{2}-\sigma(y_{1})\right)^{2}+\varepsilon(y_{1}),\end{array}\right.$
where we define $\sigma$ and $\varepsilon$ as follows, but the POf may be applied without this information. For all $x\in\mathbb{R}$, define $\sigma(x)\triangleq 2\left\lfloor{x}\right\rceil$, and $\varepsilon(x)\triangleq\left\lvert{x}\right\rvert(1+\sin(\frac{2\pi}{x})^{2})^{\frac{1}{2}}+\left\lvert{\left\lfloor{x}\right\rceil}\right\rvert$ if $x\in\mathbb{R}^{*}$ and $\varepsilon(0)\triangleq 0$, where $\left\lfloor{x}\right\rceil\triangleq\left\lfloor{x}\right\rfloor$ if $x\in\mathbb{R}_{-}$ and $\left\lfloor{x}\right\rceil\triangleq\left\lceil{x}\right\rceil-1$ if $x\in\mathbb{R}_{+}^{*}$; then the unique minimizer of $\varphi$ is $y^{*}\triangleq(0,0)$, with $\varphi(y^{*})=0$. In this context, we define $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, so the unique minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
$\gamma(x)\triangleq(x,\sigma(x)),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x),$
pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{ref}}$) minimizes
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right.$
shown on Figure 3 (right). Its global minimizer is $x^{*}\triangleq 0$, with $\Phi(x^{*})=0$, and $\Phi$ is left-discontinuous at $x^{*}$. Assumptions 1 and 2 hold. Table 1 shows that the cDSM approaches $x^{*}$ closely and from the side avoiding the discontinuity. All instances return $\widetilde{x}^{*}\in[0,2E^{-10}]$. Hence they all provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})$ such that $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 2E^{-10}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 2E^{-10}$. This empirically validates Theorem 2. We also remark that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), even if the requirement in Theorem 2 that $\varphi$ is lower semicontinuous at all points in $\mathcal{ACC}(\gamma;x^{*})$ does not hold (at $(0,-1)\in\mathcal{ACC}(\gamma;x^{*})$).
Figure 2: function $\varphi$ in Section 4.1, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and locus of $\gamma$. For each $y_{1}\in\mathbb{R}$, we plot $\varphi_{|\{y_{1}\}\times\mathbb{R}}$ for the values $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$ only.
Figure 3: functions $\varepsilon$ and $\Phi$ in Section 4.1, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
$+9.753$ $\hat{x}^{23}=+3.37E^{-03}$ $\hat{x}^{37}=+8.86E^{-07}$ $\hat{x}^{55}=+3.15E^{-09}$ $\hat{x}^{64}=+7.41E^{-12}$
$+\pi$ $\hat{x}^{18}=+3.12E^{-03}$ $\hat{x}^{31}=+3.93E^{-06}$ $\hat{x}^{49}=+1.00E^{-09}$ $\hat{x}^{56}=+6.89E^{-11}$
$+\sqrt{2}$ $\hat{x}^{13}=+4.65E^{-03}$ $\hat{x}^{29}=+2.79E^{-06}$ $\hat{x}^{44}=+3.08E^{-09}$ $\hat{x}^{54}=+5.12E^{-11}$
$+e+1$ $\hat{x}^{17}=+2.08E^{-03}$ $\hat{x}^{24}=+3.01E^{-06}$ $\hat{x}^{45}=+3.29E^{-09}$ $\hat{x}^{53}=+3.45E^{-11}$
$-9.753$ $\hat{x}^{17}=+4.61E^{-04}$ $\hat{x}^{33}=+3.33E^{-06}$ $\hat{x}^{51}=+3.96E^{-09}$ $\hat{x}^{60}=+4.05E^{-12}$
$-\pi$ $\hat{x}^{20}=+2.94E^{-03}$ $\hat{x}^{33}=+3.46E^{-06}$ $\hat{x}^{43}=+2.33E^{-10}$ $\hat{x}^{54}=+1.16E^{-10}$
$-\sqrt{2}$ $\hat{x}^{06}=+3.45E^{-03}$ $\hat{x}^{27}=+1.13E^{-06}$ $\hat{x}^{43}=+1.51E^{-09}$ $\hat{x}^{52}=+7.67E^{-13}$
$-e-1$ $\hat{x}^{21}=+2.74E^{-03}$ $\hat{x}^{36}=+2.40E^{-06}$ $\hat{x}^{49}=+4.02E^{-09}$ $\hat{x}^{57}=+5.70E^{-11}$
Table 1: results for eight instances of the cDSM in the context of Section 4.1. In this table, $\hat{x}^{k}\triangleq x^{k}$.
4.2 Radial noise
Consider an unconstrained ($\Omega\triangleq\mathbb{Y}$) case in polar coordinates ($\mathbb{Y}=Y\triangleq\mathbb{R}_{+}\times[0,2\pi[$), involving
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ (r,\theta)&\mapsto&\sqrt{r}\sin\left(\dfrac{\theta-\sigma(r)}{2}\right)^{2}+\varepsilon(r),\end{array}\right.$
where the blackboxes $\sigma$ and $\varepsilon$ are defined as follows (but the POf has no access to this information). For all $x\in\mathbb{R}_{+}$, define $\sigma(x)\triangleq\pi-2\pi\log_{2}(x)$ if $x>0$ and $\sigma(0)\triangleq 0$, and $\varepsilon(x)\triangleq\sqrt{\lvert{x^{2}-2}\rvert}+\frac{\sin(10\pi(x-\sqrt{2}))^{2}}{10}$; then the global minimizer of $\varphi$ is $y^{*}\triangleq(\sqrt{2},0)$, with $\varphi(y^{*})=0$. We define $\mathbb{Y}(x)\triangleq\{x\}\times[0,2\pi[$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}_{+}$, so the unique global minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
$\gamma(x)\triangleq(x,\sigma(x)\sim 2\pi),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x),$
where $\sigma(x)\sim 2\pi$ denotes the residual of $\sigma(x)$ modulo $2\pi$. Figure 4 (left) draws $\varphi$, Figure 4 (right) shows the locus of $\gamma$, and Figure 5 (left) pictures $\varepsilon$. Then the function $\Phi$ equals
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right.$
shown on Figure 5 (right). Its minimizer is $x^{*}\triangleq\sqrt{2}$, with $\Phi(x^{*})=0$. Assumptions 1 and 2 hold. Table 2 shows the convergence of the cDSM towards $x^{*}$. All instances return $\widetilde{x}^{*}\in[x^{*}\pm 6E^{-11}]$, so $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})$ satisfies $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 6E^{-11}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 3E^{-5}$. This agrees with Theorem 2.
Figure 4: function $\varphi$ in Section 4.2, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and locus of $\gamma$.
Figure 5: functions $\varepsilon$ and $\Phi$ in Section 4.2, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
00 $\hat{x}^{09}=+4.39E^{-03}$ $\hat{x}^{29}=-8.47E^{-07}$ $\hat{x}^{41}=+2.73E^{-09}$ $\hat{x}^{52}=+5.39E^{-11}$
$2^{-5}$ $\hat{x}^{09}=+5.08E^{-04}$ $\hat{x}^{27}=+4.42E^{-06}$ $\hat{x}^{44}=-3.57E^{-09}$ $\hat{x}^{54}=+3.83E^{-11}$
$3\sqrt{2}$ $\hat{x}^{17}=-1.14E^{-03}$ $\hat{x}^{35}=-1.51E^{-06}$ $\hat{x}^{51}=-1.95E^{-09}$ $\hat{x}^{60}=+2.94E^{-11}$
$4\pi$ $\hat{x}^{25}=+1.37E^{-03}$ $\hat{x}^{39}=+1.50E^{-06}$ $\hat{x}^{57}=+2.05E^{-09}$ $\hat{x}^{66}=-4.99E^{-11}$
$5$ $\hat{x}^{08}=+6.72E^{-05}$ $\hat{x}^{27}=-1.45E^{-06}$ $\hat{x}^{41}=-1.81E^{-09}$ $\hat{x}^{48}=+5.71E^{-11}$
$e$ $\hat{x}^{08}=-1.75E^{-03}$ $\hat{x}^{28}=+9.13E^{-07}$ $\hat{x}^{41}=+4.04E^{-09}$ $\hat{x}^{53}=-3.28E^{-11}$
$e^{2}$ $\hat{x}^{11}=-5.45E^{-04}$ $\hat{x}^{25}=+4.01E^{-06}$ $\hat{x}^{43}=+5.27E^{-10}$ $\hat{x}^{53}=-5.47E^{-11}$
$e^{3}$ $\hat{x}^{30}=+3.83E^{-03}$ $\hat{x}^{44}=+4.66E^{-06}$ $\hat{x}^{59}=-4.19E^{-09}$ $\hat{x}^{69}=+3.72E^{-12}$
Table 2: results for eight instances of the cDSM in the context of Section 4.2. In this table, $\hat{x}^{k}\triangleq x^{k}-\sqrt{2}$.
4.3 Noise affected by a nonlinear combination of all variables
Consider the variables space $\mathbb{Y}\triangleq\mathbb{R}^{2}$, the feasible set $\Omega\triangleq\mathbb{R}_{+}\times\mathbb{R}$ and the function
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\overline{\mathbb{R}}\\ y\triangleq(y_{1},y_{2})&\mapsto&\ln\left(1+\left(\dfrac{y_{1}^{2}}{y_{2}^{2}+1}-1\right)^{2}\right)+\varepsilon(f(y)),\end{array}\right.$
where $f:y\in\mathbb{Y}\mapsto y_{1}y_{2}\in\mathbb{R}$ is assumed explicit, while $\varepsilon$ is defined via $\varepsilon(x)\triangleq\exp(\frac{1}{x-4})+\frac{\sqrt{\lvert{x-4}\rvert}}{5}$ if $x\neq 4$ and $\varepsilon(4)\triangleq+\infty$ and this information is assumed hidden. Then $\varphi$ has no global minimizer, but there exists a generalized global solution $y^{*}\approx(2.1287,1.8791)$, since $\inf\varphi(\Omega)=0$ and $\varphi(y)\to 0$ when $y\to y^{*}$ from some direction, even though $\varphi(y^{*})=+\infty$. The exact value of $y^{*}$ and the direction are given below. We define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, so $\varphi_{|\mathbb{Y}(x)}$ is smooth (since $f(y)=x$ so $\varepsilon(f(y))$ is constant) and its minimizer is computable: First, $y_{1}^{2}/(y_{2}^{2}+1)=1$ provides $y_{2}^{2}=y_{1}^{2}-1$ (and thus $y_{1}^{2}\geq 1$ so $y_{1}\neq 0$), and $(y_{1},y_{2})\in\mathbb{Y}(x)\cap\Omega$ raises $y_{2}=x/y_{1}$ and $y_{1}\geq 0$. Second, equalling the two expressions of $y_{2}^{2}$ provides $(y_{1}^{2})^{2}-y_{1}^{2}-x^{2}=0$, with one admissible value for $y_{1}^{2}$ as $y_{1}^{2}=\frac{1}{2}(1+\sqrt{1+4x^{2}})$ since $y_{1}^{2}\geq 1$. Hence $y_{1}$ follows (since $y\in\Omega$ forces $y_{1}\geq 0$), and then $y_{2}$ follows as well. Then, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
$\gamma(x)\triangleq\left(\sqrt{\frac{1+\sqrt{1+4x^{2}}}{2}},x\sqrt{\frac{2}{1+\sqrt{1+4x^{2}}}}\right),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x).$
Figure 6 shows $\varphi$, the partition of $\mathbb{Y}$ and the locus of $\gamma$, and Figure 7 (left) shows $\varepsilon$. Remark that $\gamma$ is continuous, even if $\varphi$ is not. In this context, $\Phi$ equals
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right.$
shown on Figure 7 (right). Then $\Phi$ has no minimum but $x^{*}\triangleq 4$ is a generalized global minimizer, since $\lim_{x\nearrow x^{*}}\Phi(x)=0=\inf\Phi(\mathbb{X})$ despite $\lim_{x\searrow x^{*}}\Phi(x)=+\infty=\Phi(x^{*})$. Thus, $y^{*}\triangleq\gamma(x^{*})$ is a generalized global minimizer of $\varphi$ and $\lim_{y=\gamma(x\nearrow x^{*})}\varphi(y)=0$. Assumptions 1 and 2 hold.
We test eight instances of the cDSM minimizing $\Phi$. Table 3 shows their convergence towards $x^{*}$. All instances starting with $x^{0}<x^{*}$ return $\widetilde{x}^{*}\in[x^{*}-2E^{-10},x^{*}]$, which approaches $x^{*}$ closely and from the direction minimizing $\Phi$. Hence they provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})\approx y^{*}$ with $\varphi(\widetilde{y}^{*})\approx\inf\varphi(Y)$, which agrees with Theorem 2. The instance starting with $x^{0}\triangleq 3\sqrt{2}\approx 4.24>x^{*}$ behaves similarly. The last three instances, which start with $x^{0}\gg 4$, converge instead to $\widetilde{x}^{*}\approx 9.27$, but this is consistent with Theorem 2 since $\Phi$ has a local minimizer around $9.27$ so $\varphi$ has a local minimizer around $\gamma(9.27)$.
Figure 6: function $\varphi$ in Section 4.3, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and the locus of $\gamma$.
Figure 7: functions $\varepsilon$ and $\Phi$ in Section 4.3, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$. Note that neither $\varepsilon$ nor $\Phi$ have a global minimum.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
$-e^{2}$ $\hat{x}^{107}=-4.17E^{-03}$ $\hat{x}^{120}=-2.20E^{-06}$ $\hat{x}^{138}=-1.68E^{-09}$ $\hat{x}^{146}=-4.65E^{-11}$
$-\pi$ $\hat{x}^{026}=-1.76E^{-03}$ $\hat{x}^{043}=-1.70E^{-06}$ $\hat{x}^{057}=-3.88E^{-09}$ $\hat{x}^{065}=-4.14E^{-11}$
$-\sqrt{2}$ $\hat{x}^{024}=-4.25E^{-04}$ $\hat{x}^{040}=-4.95E^{-06}$ $\hat{x}^{051}=-1.40E^{-09}$ $\hat{x}^{062}=-8.01E^{-12}$
$+e$ $\hat{x}^{009}=-4.68E^{-04}$ $\hat{x}^{026}=-2.78E^{-06}$ $\hat{x}^{043}=-3.13E^{-09}$ $\hat{x}^{051}=-1.06E^{-10}$
$+3\sqrt{2}$ $\hat{x}^{017}=-2.66E^{-03}$ $\hat{x}^{030}=-1.33E^{-07}$ $\hat{x}^{046}=-2.41E^{-09}$ $\hat{x}^{053}=-8.40E^{-11}$
$+2e^{2}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
$+4\pi$ / / / $\hat{x}^{201}=+5.27E^{+00}$
$+e^{3}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
Table 3: results for eight instances of Algorithm 1 in the context of Section 4.3. In this table, $\hat{x}^{k}\triangleq x^{k}-4$. For each of the last three instances, we arbitrarily return $x^{201}\approx 9.26779505$ since $(x^{k})_{k\geq 40}$ has its first $9$ digits constant in all cases.
4.4 Bidimensional noise and oracle provided by an algorithmic method
Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{3}$, and $\Omega\triangleq\mathbb{Y}$, and the function
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ y&\mapsto&\left\lVert{y}\right\rVert_{\infty}+\varepsilon\left(f(y)\right),\end{array}\right.$
where $f$, defined by $f(y)\triangleq(y_{2}-y_{1}^{3},y_{1}-y_{3}^{3})$ for all $y\triangleq(y_{1},y_{2},y_{3})\in\mathbb{Y}$, is explicitly known while the blackbox $\varepsilon$ is defined by $\varepsilon(x)\triangleq(\frac{\sin(10\pi(x_{2}-x_{1}^{3}))}{5}+\frac{\sin(6\pi(x_{2}-e^{-x_{1}}+1))}{7}+\frac{\sin(12\pi(x_{1}^{2}+x_{2}^{2})^{\frac{1}{2}})}{11})^{2}$ for all $x\in\mathbb{R}^{2}$. The global minimizer of $\varphi$ is $y^{*}\triangleq(0,0,0)$, with $\varphi(y^{*})=0$. In this context, we define the partition sets of $\mathbb{Y}$ as $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}^{2}$, so $\varepsilon_{|\mathbb{Y}(x)}$ is constant. For all $x\triangleq(x_{1},x_{2})\in\mathbb{X}$, some calculation allow us to rewrite $\mathbb{Y}(x)=\{(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}}),\leavevmode\nobreak\ t\in\mathbb{R}\}$, so the unique minimizer of $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varepsilon(x)+\max\left\{\left\lvert{t_{x}(y)}\right\rvert,\left\lvert{t_{x}(y)^{3}+x_{1}}\right\rvert,\left\lvert{\sqrt[3]{t_{x}(y)-x_{2}}}\right\rvert\right\},$
where, for any $y\in\mathbb{Y}(x)$$t_{x}(y)$ denotes the unique $t\in\mathbb{R}$ such that $s_{x}(t)\triangleq(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}})=y$. This solution is hardly tractable. However, the problem admits a smooth exact reformulation given by
$\begin{array}[]{cl}\underset{\begin{array}[]{c}M\geq 0,\leavevmode\nobreak\ T\in\mathbb{R}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\varepsilon(x)+M\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}T\in I_{x}^{1}(M)\triangleq\left[-M,M\right],\\ T\in I_{x}^{2}(M)\triangleq\left[\sqrt[3]{-M-x_{1}},\sqrt[3]{M-x_{1}}\right],\\ T\in I_{x}^{3}(M)\triangleq\left[-M^{3}+x_{2},M^{3}+x_{2}\right].\end{array}\end{array}$
Its unique solution, denoted by $(M(x),T(x))$, provides
$\gamma(x)\triangleq s_{x}(T(x)),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x)+M(x).$
Figure 8 shows $\varphi$, Figure 9 draws the partition of $\mathbb{Y}$ and the locus of $\gamma$, and Figure 10 (left) and (center) represent respectively $\varepsilon$ and $M$. In this context, $\Phi$ equals
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x)+M(x),\end{array}\right.$
pictured on Figure 10 (right). Its minimizer is $x^{*}\triangleq(0,0)$, with $\Phi(x^{*})=0$. Assumptions 1 and 2 hold.
Note that $(M(x),T(x))$ remains intractable analytically. Nevertheless, we observe that
$\begin{array}[]{ccl}M(x)&=&\min\left\{M\in\mathbb{R}_{+}\ :\ I_{x}(M)\triangleq I_{x}^{1}(M)\cap I_{x}^{2}(M)\cap I_{x}^{3}(M)\neq\emptyset\right\},\\ \left\{T(x)\right\}&=&I_{x}\left(M(x)\right),\end{array}$
so $M(x)$ is obtainable via a dichotomic search and the singleton $\{T(x)\}$ follows. In practice, running sufficiently many iterations of the dichotomic search provides an accurate approximation of $M(x)$. We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $[M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$. This allows to define $M(x)\approx\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})$, and then $T(x)\approx\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))$, and finally $\gamma(x)\approx\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))$.
A total of eight instances of the cDSM are tested. They actually minimize $\hat{\Phi}\triangleq\varepsilon+\hat{M}$ since we approximate $M\approx\hat{M}$ for tractability. Table 4 shows the convergence of these instances towards $x^{*}$. They all return $\left\lVert{\widetilde{x}^{*}-x^{*}}\right\rVert\leq 9E^{-7}$, hence they provide $\widetilde{y}^{*}\triangleq\hat{\gamma}(\widetilde{x}^{*})$ such that $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 9E^{-7}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 9E^{-7}$. Five instances even return $\left\lVert{\widetilde{x}^{*}-x^{*}}\right\rVert\leq 5E^{-10}$. These behaviours agree with Theorem 2, even if the problem we actually solve slightly differs from the true problem.
Figure 8: function $\varphi$ in Section 4.4, restriction over planes with constant third variable.
Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section 4.4.
Figure 10: functions (left) $\varepsilon$, (center) $M$ and (right) $\Phi$ in Section 4.4. In this section $\Phi=\varepsilon+M$.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-02}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-05}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-08}$ returned $\hat{x}^{k}$
$(-2,2)$ $\hat{x}^{032}=3.52E^{-02}$ $\hat{x}^{110}=4.03E^{-05}$ $\hat{x}^{194}=4.79E^{-08}$ $\hat{x}^{220}=1.10E^{-08}$
$(\frac{-1}{100},e^{2})$ $\hat{x}^{037}=2.02E^{-02}$ $\hat{x}^{100}=3.99E^{-05}$ $\hat{x}^{176}=1.10E^{-08}$ $\hat{x}^{198}=9.03E^{-10}$
$(\frac{-\pi}{2},\frac{7}{4})$ $\hat{x}^{027}=6.34E^{-03}$ $\hat{x}^{095}=4.60E^{-05}$ $\hat{x}^{177}=1.23E^{-08}$ $\hat{x}^{195}=5.51E^{-09}$
$(\frac{-\pi}{4},e^{\frac{1}{2}})$ $\hat{x}^{040}=3.66E^{-02}$ $\hat{x}^{126}=1.46E^{-05}$ $\hat{x}^{235}=3.45E^{-08}$ $\hat{x}^{253}=4.54E^{-09}$
$(\frac{1}{4},\frac{1}{4})$ $\hat{x}^{022}=2.85E^{-02}$ $\hat{x}^{119}=2.90E^{-05}$ $\hat{x}^{193}=1.84E^{-08}$ $\hat{x}^{212}=2.88E^{-09}$
$(\frac{3\pi}{2},\frac{1}{\sqrt{8}})$ $\hat{x}^{039}=3.29E^{-02}$ $\hat{x}^{124}=1.96E^{-05}$ $\hat{x}^{186}=4.86E^{-08}$ $\hat{x}^{226}=5.22E^{-09}$
$(e^{2},2\pi)$ $\hat{x}^{036}=3.43E^{-02}$ $\hat{x}^{109}=3.00E^{-05}$ $\hat{x}^{171}=4.40E^{-08}$ $\hat{x}^{211}=8.19E^{-09}$
$(e^{2},\frac{-1}{11})$ $\hat{x}^{029}=1.48E^{-02}$ $\hat{x}^{093}=4.77E^{-05}$ $\hat{x}^{148}=2.48E^{-08}$ $\hat{x}^{172}=1.48E^{-09}$
Table 4: results for eight instances of Algorithm 1 in the context of Section 4.4. In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
5 Numerical gain of performance provided by the POf
In this section, we consider again composite problem as in Section 4, where the POf allows to focus on Reformulation ($\mathbf{P}_{\texttt{ref}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) instead of Problem ($\mathbf{P}_{\texttt{ini}}$) with $\dim(\mathbb{Y})$ possibly large. We alter the problems from Section 4 so that $\dim(\mathbb{Y})\approx 100$ while $\dim(\mathbb{X})\leq 10$. Let us stress that our choice to consider $\dim(\mathbb{Y})$ at most $100$ stands only for ease of presentation and limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [10]).
In most examples below, computing $\varphi(y)$$\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$. This likely differs from real cases, where either $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by solving an optimization subproblem. Thus, to establish a realistic comparison between the solvers, we proceed as follows. For each problem, we consider that the cost to compute $\varphi(y)$ is $1$ unit, identical for all $y\in\mathbb{Y}$, and that the computation of $\Phi(x)$ costs $1+\tau$ units, identical for all $x\in\mathbb{X}$ and where $\tau\geq 0$ is fixed. The value $\tau$ captures the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$, so the computational cost of $\Phi(x)$ models the additional cost required to first compute $\gamma(x)$ before computing $\varphi(\gamma(x))$. Then, for each method, we track the sequence of all evaluated points. This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{ini}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{ref}}$) instead. Finally, for each method we plot the graph representing the lowest objective value found within a budget of $c$ units, for all $c\in\mathbb{R}_{+}$.
We arbitrarily choose as baseline two popular DFO solvers, NOMAD [6] and PRIMA [30], solving Problem ($\mathbf{P}_{\texttt{ini}}$). We compare them to our strategy formalized in Theorem 2. This shows that if $\tau$ is not prohibitively large, then our method significantly outperforms both NOMAD and PRIMA. Note that this observation does not discredit these solvers: they are not designed for such large-dimensional contexts, and more suited DFO techniques [19, 21, 22, 23, 26] may also perform poorly when we rise $\dim(\mathbb{Y})$ sufficiently high (up to $\dim(\mathbb{Y})=+\infty$ as in [10]). Similarly, our results do not give credit to our naive implementation of the cDSM. What makes our methodology efficient is the sheer dimension reduction provided by the POf, and the guarantee of (possibly, generalized) local optimality of points returned by any algorithm with a covering step.
We follow the extreme barrier approach [3] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$ instead of $\varphi$ over $\Omega$, and we provide relevant box constraints to NOMAD and PRIMA. We run NOMAD with default parameters, PRIMA with the BObyQA algorithm (to handle the box constraints), and the cDSM as in Remark 1. We give to NOMAD and PRIMA six random starting points $(y^{0}_{\ell})_{\ell\in\llbracket 1,6\rrbracket}$, and we give to our cDSM the starting points $(\chi(y^{0}_{\ell}))_{\ell\in\llbracket 1,6\rrbracket}$. In this section, we denote by $\mathds{1}$ the vector with all entries equal to $1$.
5.1 101-dimensional problem with mono-variable noise
Consider an alteration of the example from Section 4.1 given by $\mathbb{Y}=Y\triangleq\mathbb{R}^{101}$, and $\Omega\triangleq\mathbb{Y}$, and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ y\triangleq(y_{i})_{i=0}^{100}&\mapsto&\left\lVert{y-\sigma(y_{0})}\right\rVert^{2}+\varepsilon(y_{0}),\end{array}\right.$
where $\varepsilon(x)\triangleq\left\lvert{x}\right\rvert(1+\sin(\frac{2\pi}{x})^{2})^{\frac{1}{2}}+\left\lvert{\left\lfloor{x}\right\rceil}\right\rvert$ if $x\in\mathbb{R}^{*}$ and $\varepsilon(0)\triangleq 0$ as in Section 4.1, and where
$\sigma:\left\{\begin{array}[]{ccl}\mathbb{R}&\to&\mathbb{R}^{101}\\ x&\mapsto&\left(\leavevmode\nobreak\ x\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(2\left(1+\frac{i-1}{5}\right)\left\lfloor{\frac{x}{i}}\right\rceil\right)_{i=1}^{25}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(25\sin\left(\frac{i-25}{5}\pi x\right)\right)_{i=26}^{50}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(x-\frac{10}{i}\right)_{i=51}^{75}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(\frac{i}{10}\right)_{i=76}^{100}\leavevmode\nobreak\ \right).\end{array}\right.$
We consider $\sigma$ and $\varepsilon$ as blackboxes so their expression is assumed to be unknown to all methods, but we remark that the global minimizer of $\varphi$ is $y^{*}\triangleq\sigma(0)$, with $\varphi(y^{*})=0$. We define $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}^{100}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, since therefore $\varphi_{|\mathbb{Y}(x)}$ admits an easily tractable global minimizer as
$\gamma(x)\triangleq\sigma(x),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(y_{0}).$
Thus, Assumptions 1 and 2 hold and $\Phi=\varepsilon$ is as in Section 4.1. Figure 11 compares the three methods.
Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $1$-dimensional DFO problem that focuses on $y_{0}$, and the POf highlights how to handle the $100$ others in the original problem. Even by considering $\tau=100$, the POf provides a sensible gain of performance: our naive cDSM relying on the POf outperforms the solvers solving the original problem. Its returned solution has a significantly lower objective value; and the computational budget is comparable if the pointwise cost to evaluate $\gamma$ is $100$ times greater than those of $\varphi$. Our cDSM converges towards the global minimizer (the returned objective value is around $1E^{-8}$ is all cases), while NOMAD and PRIMA solving the original problem remain both in an exploratory phase that returns poor results. The early stop of PRIMA in the first case even shows that PRIMA interrupted itself on a poor incumbent solution with $40\%$ of its budget remaining.
Figure 11: Comparison of the best solution found by each solver depending on the computational budget spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[-30,30]^{101}$.
5.2 Radial noise in a 101-dimensional problem
Consider an alteration of the problem from Section 4.2, with $\mathbb{Y}=Y=\Omega\triangleq\mathbb{R}_{+}\times[0,2\pi[^{100}$ and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ \left(r,(\theta_{i})_{i=1}^{100}\right)&\mapsto&\dfrac{\sqrt{r}}{100}\sum\limits_{i=1}^{100}\sin\left(\dfrac{\theta_{i}-\sigma_{i}(r)}{2}\right)^{2}+\varepsilon(r),\end{array}\right.$
where $\sigma(x)\triangleq(2\pi\log_{i+1}(x))_{i=1}^{100}$ for all $x>0$ and $\sigma(0)\triangleq 0\mathds{1}$, and $\varepsilon(x)\triangleq\sqrt{\lvert{x^{2}-2}\rvert}+\frac{\sin(10\pi(x-\sqrt{2}))^{2}}{10}$ for all $x\in\mathbb{R}_{+}$, are two blackboxes. The global minimizer of $\varphi$ is $y^{*}\triangleq(\sqrt{2},\sigma(\sqrt{2}))$, with $\varphi(y^{*})=0$. We define $\mathbb{Y}(x)\triangleq\{x\}\times[0,2\pi[^{100}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}_{+}$. Then, $\varphi_{|\mathbb{Y}(x)}$ is a smooth function of $\theta$ while $r=x$ is fixed, and its global minimizer is
$\gamma(x)\triangleq(x,\sigma(x)\sim 2\pi),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x),$
where $\sigma(x)\sim 2\pi$ denotes the component-wise residual of $\sigma(x)$ modulo $2\pi$. Thus, Assumptions 1 and 2 hold and $\Phi=\varepsilon$ is as in Section 4.2. The comparison between the strategies is provided in Figure 12.
The results are similar to those in Section 5.1. Even by considering $\tau=100$, the POf provides a gain of performance. Within a comparable budget, our cDSM converges towards the global minimizer, while NOMAD and PRIMA solving the original problem remain in an exploratory phase that returns irrelevant points. The very early stop of NOMAD in three cases even shows that NOMAD sometimes interrupts itself early on its poor incumbent solution.
Figure 12: Comparison of the best solution found by each solver depending on the computational cost spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[0,2\pi]^{101}$.
5.3 Noise affected by a nonlinear combinations of all 100 variables
Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{100}$, and $\Omega\triangleq\{y\in\mathbb{Y}:\forall i\in\llbracket 1,100\rrbracket,\leavevmode\nobreak\ y_{i+1}\geq y_{i}>0\}$, and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ y\triangleq(y_{i})_{i=1}^{100}&\mapsto&\sum\limits_{(\ell,p)\in\llbracket 0,19\rrbracket^{2}}\ln\left(1+\left(\dfrac{\pi_{\ell}(y)}{\pi_{p}(y)}-1\right)^{2}\right)+\varepsilon\left(f(y)\right),\end{array}\right.$
where for all $y\in\mathbb{Y}$$\pi_{\ell}(y)\triangleq\prod_{i=5\ell+1}^{5(\ell+1)}y_{i}$ for each $\ell\in\llbracket 0,19\rrbracket$ and $f(y)\triangleq\frac{1}{5}\sum_{\ell=0}^{19}\pi_{\ell}(y)$ have explicit definition, and where the blackbox $\varepsilon$ satisfies $\varepsilon(x)\triangleq\exp(\frac{1}{z-4})+\frac{1}{5}\sqrt{\lvert{z-4}\rvert}$ if $z\neq 4$ and $\varepsilon(4)\triangleq+\infty$. Here $\varphi$ has no global minimizer, but $\inf\varphi(\Omega)=0$ and $y^{*}\triangleq\mathds{1}$ is a generalized global solution (even though $\varphi(y^{*})=+\infty$). For all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, we define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$. Then, $\varphi_{|\mathbb{Y}(x)}$ is a smooth function since $\varepsilon(f(y))$ is constant. Moreover its minimizer is analytically obtainable as follows. For all $x\in\mathbb{X}$$\varphi_{|\mathbb{Y}(x)}(y)\geq 0$ and equals 00 when all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ are equal. Under the constraint $y\in\Omega$, equalling all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ implies that $y=\alpha\mathds{1}$ for some $\alpha\in\mathbb{R}$. Then, $y\in\mathbb{Y}(x)$ leads to $4\alpha^{5}=x$. The value of $\alpha$ follows and we get that
$\gamma(x)\triangleq\sqrt[5]{\frac{x}{4}}\mathds{1},\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x).$
Thus, $\Phi=\varepsilon$ as in Section 4.3, and Assumptions 1 and 2 hold. Figure 13 compares our strategies.
Similarly to Sections 5.1 and 5.2, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly. Even with $\tau=100$, the POf provides a gain of performance. NOMAD and PRIMA solving the original problem both fail to significantly improve their initial incumbent solution and, in the second case, both solvers behave as if their initial incumbent solution cannot be improved. Within a comparable budget, our cDSM converges towards the global minimizer on all but one case. Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by NOMAD and PRIMA, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{ref}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large. Our instance lacks one since it lacks a search step.
Figure 13: Comparison of the best solution found by each solver depending on the computational cost spent. The starting points are chosen as $y^{0}_{1}\triangleq(i/100)_{i=1}^{100}$$y^{0}_{2}\triangleq 0.5\mathds{1}$$y^{0}_{3}\triangleq(\frac{2i}{100})_{i=1}^{100}$, and $y^{0}_{i}$, for each $i\in\llbracket 4,6\rrbracket$, is chosen randomly but such that for each $j\in\llbracket 1,100\rrbracket$, the $j^{th}$ component of $y^{0}_{i}$ lies in $[\frac{j-1}{100},\frac{j}{100}]$.
5.4 Ten-dimensional noise with non-analytical oracle and 100 variables
Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{100}$, and $\Omega\triangleq\mathbb{Y}$, and the function
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{R}^{100}&\to&\mathbb{R}\\ y&\mapsto&\left\lVert{y}\right\rVert_{1}+\varepsilon(f(y)),\end{array}\right.$
where the function $f$ (with definition assumed accessible) is defined as follows, with $\mathbb{X}=X\triangleq\mathbb{R}^{10}$,
$f:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{X}\\ (y_{i})_{i=1}^{100}&\mapsto&\left(g_{1}(y_{10})-\sum\limits_{i=1}^{9}y_{i},\leavevmode\nobreak\ \dots,\leavevmode\nobreak\ g_{10}(y_{100})-\sum\limits_{i=91}^{99}y_{i}\right),\end{array}\right.$
with $g_{j}(z)\triangleq(z+(1+\frac{j}{10})^{z}-1)$ for all $z\in\mathbb{R}$ and all $j\in\mathbb{N}$, and where the blackbox $\varepsilon$ is defined through
$\varepsilon:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ z=(z_{j})_{j=1}^{10}&\mapsto&\left(\frac{\sin 5\pi\left(z_{2}-z_{1}^{3}\right)}{5}+\frac{\sin 6\pi\left(z_{4}-e^{-z_{2}-z_{3}}+1\right)}{7}+\frac{\sin 7\pi\sqrt{z_{5}^{2}+z_{6}^{2}+z_{7}^{2}}}{11}+\frac{\sin 8\pi z_{8}z_{9}z_{10}}{13}\right)^{2}.\end{array}\right.$
The global minimizer of $\varphi$ is $y^{*}\triangleq 0\mathds{1}$, with $\varphi(y^{*})=0$. We partition $\mathbb{Y}$ via $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$. Then, for all $x\in\mathbb{X}$, the constraint $y\in\mathbb{Y}(x)$ fixes all variables $(y_{10j})_{j=1}^{10}$ to $g_{j}^{-1}(x_{j}+\sum_{i=10(j-1)+1}^{10(j-1)+9}y_{i})$, and $\varepsilon(f(y))$ to the constant $\varepsilon(x)$, so $\gamma(x)$ minimizing $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varepsilon(x)+\sum\limits_{j=1}^{10}\left(\sum\limits_{i=10(j-1)+1}^{10(j-1)+9}\left\lvert{y_{i}}\right\rvert+\left\lvert{g_{j}^{-1}\left(x_{j}+\sum\limits_{i=10(j-1)+1}^{10(j-1)+9}y_{i}\right)}\right\rvert\right).$
Then, observing that $\frac{\mathrm{d}g_{j}^{-1}}{\mathrm{d}z}(z)\in{]}0,1{[}$ for all $(j,z)\in\mathbb{N}^{*}\times\mathbb{R}$, a sensitivity analysis with respect to each sum $\sum_{i=10(j-1)+1}^{10(j-1)+9}\left\lvert{y_{i}}\right\rvert$ shows that
$\gamma(x)\triangleq(\gamma_{i}(x))_{i=1}^{100}\quad\mbox{where}\quad\gamma_{i}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\leavevmode\nobreak\ i\notin 10\mathbb{N},\\ g_{j}^{-1}(x_{j})&\mbox{if}\leavevmode\nobreak\ i=10j,\leavevmode\nobreak\ j\in\llbracket 1,10\rrbracket,\end{array}\right.$
with $\varphi(\gamma(x))=\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{g_{j}^{-1}(x_{j})}\right\rvert$. As a result, $\Phi$ equals
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x\triangleq(x_{j})_{j=1}^{10}&\mapsto&\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{g_{j}^{-1}(x_{j})}\right\rvert.\end{array}\right.$
Its global minimizer is $x^{*}\triangleq 0\mathds{1}$, with $\Phi(x^{*})=0$. Assumptions 1 and 2 hold.
Unfortunately, $\gamma$ admits no analytical expression since, for all $j\in\llbracket 1,10\rrbracket$$g_{j}^{-1}$ cannot be expressed with elementary functions. However, since $g_{j}^{-1}(x_{j})$ solves the equation $(z+(1+\frac{j}{10})^{z}-1)-x_{j}=0$ (with variable $z\in\mathbb{R}$), we approximately solve this equation using a dichotomic search to obtain an approximate solution $\hat{g_{j}}^{-1}(x_{j})\approx g_{j}^{-1}(x_{j})$. As a result, for all $x\in X$, we approximate
$\gamma(x)=(\gamma_{i}(x))_{i=1}^{100}\approx\hat{\gamma}(x)\triangleq(\hat{\gamma_{i}}(x))_{i=1}^{100}\quad\mbox{where}\quad\hat{\gamma_{i}}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\leavevmode\nobreak\ i\notin 10\mathbb{N},\\ \hat{g_{j}}^{-1}(x_{j})&\mbox{if}\leavevmode\nobreak\ i=10j,\leavevmode\nobreak\ j\in\llbracket 1,10\rrbracket,\end{array}\right.$
and
$\Phi\approx\hat{\Phi}\quad\mbox{where}\quad\hat{\Phi}:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x\triangleq(x_{j})_{j=1}^{10}&\mapsto&\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{\hat{g_{j}}^{-1}(x_{j})}\right\rvert.\end{array}\right.$
Figure 14 shows a comparison between the different strategies, where for tractability our naive cDSM actually minimizes $\hat{\Phi}$ instead of $\Phi$. Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 5.15.2 and 5.3. Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of the original problem accordingly. Using our cDSM, the POf provides a gain of performance if we allow $\tau\approx 10$ at most. With $\tau=10$, the POf allows our cDSM to converge under a computational budget similar to whose required by NOMAD and PRIMA, but the solutions returned by our cDSM are $21E$ 2 times better than those returned by NOMAD and $71E$ 7 times better than those returned by PRIMA. The value $\tau=10$ is quite low, but we could consider a higher value by using a better implementation of our algorithm. For example, a test with NOMAD solving the reformulated problem (instead of our naive instance of cDSM) provided similar graphs with $\tau=30$. This last run is represented on Figure 15.
Figure 14: Comparison of the best solution found by each solver depending on the computational cost spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[-5,5]^{100}$.
Figure 15: Additional gain provided by solving the reformulated problem using a more efficient solver. Despite the higher value for $\tau$ compared to the results on Figure 14, the graphs are similar.
6 General discussion
We conclude this work with a discussion on the POf. Section 6.1 summarizes the most important aspects of the POf, and Section 6.2 lists some routes for improvements.
6.1 Comments on the POf
The most stringent aspect of the POf is the oracle function $\gamma$. For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) must have a computable global solution. Ensuring this requirement is challenging in general and likely requires a problem-dependent strategy. Nevertheless, Sections 4.4 and 5.4 show that this theoretical requirement may be slightly relaxed in practice. The POf does not need that $\gamma$ has a closed-form expression, as $\gamma$ may be defined pointwise as, for all $x\in X$, the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with great confidence to approximate closely a global solution. This remains restrictive, but some applications already fit in this framework (as discussed in Sections 1.24 and 5), and we are working on possible ways to relax this requirement (see Section 6.2).
Our assumptions are another nontrivial aspect of the POf. Requirements about $\chi$ are presumably easy to satisfy since $\chi$ follows directly from the partition (which is chosen by the user), but we are not aware of any easy and exhaustive rule that ensures Assumption 1 and 2. Nevertheless, they become tractable in some cases. For example, in the context studied in Section 4, Assumptions 1 and 2 hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [2]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets.
A careful reading of the proof of Theorem 1 shows that Theorem 1.a) actually holds under no assumption except the applicability of the POf, and that Theorem 1.b) holds as soon as $\chi$ is continuous. Hence, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). The POf sometimes leads to an instance where this is possible. A nontrivial example from optimal control theory is given in [9, Chapter 7.3.1].
Theorem 2 is stronger if we moreover assume that $\varphi$ is continuous and $\Omega$ is closed. In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) for all $x^{*}$ (since $\varphi$ is therefore lower semicontinuous and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ naturally holds). In our analysis where this additional assumption does not holds, Theorem 2 states only a sufficient but not necessary condition to claim that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$). Sections 4.1 and 4.3 illustrate cases where $\varphi$ is not lower-semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed by Theorem 2.
Theorem 2 is stated for the cDSM only, but it is actually compatible with many others DFO algorithms. As discussed in Section 3.1, Theorem 2 remains valid if the cDSM in Algorithm 1 is replaced by any DFO algorithm enhanced with a covering step. In general, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is akin to a generic blackbox optimization problem, so there is no reason a priori to favour a DSM instead of another class of methods. In particular, if $\dim(\mathbb{X})$ is too large, the cDSM may not be suited to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) while another DFO may be more appropriate. We refer to [2] for a guideline showing how to indeed fit a covering into any existing DFO algorithm. However, to our best knowledge, no popular DFO solver implements some algorithms with a covering step yet. Then, in practical settings, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality). Thus Theorem 2 is not yet compatible with most DFO solvers.
The usual case of application of the POf is when $\dim(\mathbb{Y})\geq 2$ and $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$. It is advised to ensure that $\dim(\mathbb{X})\leq 50$ to avoid bad performance when solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) with the cDSM [5, Section 1.4]. Although our examples in Section 5 illustrate cases with $\dim(\mathbb{Y})=101$ at most, we stress that the performance of the POf depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$. Hence $\dim(\mathbb{Y})$ may be possibly infinite, as in [10]. Moreover, the smaller $\dim(\mathbb{X})$ is and the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is, the larger the constant $\tau$ from Section 5 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be. As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly, then the POf would be deemed useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$.
In Section 4, we show that the POf tackles a class of greybox composite problem. Although the literature on composite problem is rich, [21, 23] claim that most papers assume specific structures for $\widetilde{\varphi}$ and $\sigma$ and $\varepsilon$. We observe that our approach has complementarity with those in [21, 23]. Indeed, [21] considers cases where $\widetilde{\varphi}$ is cheap to evaluate while $\sigma$ and $\varepsilon$ are possibly expensive, but they consider $\widetilde{\varphi}$ with a complex structure and $\dim(\mathbb{Y})$ small and $\dim(\mathbb{X})\leq 65$, while we better tackle cases with $\widetilde{\varphi}$ with simple structure and $\dim(\mathbb{Y})$ large but $\dim(\mathbb{X})$ small. Also, [23] also considers $\widetilde{\varphi}$ well structured and cheap to evaluate, but unlike us it assumes that $\sigma$ and $\varepsilon$ are smooth (although expensive to calculate).
6.2 Perspectives for future work
A future work will relax the definition of $\gamma(x)$ as a tractable global solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for all $x\in X$. We plan to alter the POf to allow for the access to only an approximation of a local solution. This would be representative of most practical cases, where all subproblems are solved numerically by an optimization method which only approximates a local solution. We conducted some tests following this idea on some problems from [9, Chapter 7] by defining, for all $x\in X$$\gamma(x)$ as the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). This approach seems promising and works well in our experiments, although the theory is complex so we are still working to simplify it. This future work may also determine if, given a point $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{ref}}$), the point $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{ini}}$). This would allow for an adaptation of Theorems 1 and 2 where Reformulation ($\mathbf{P}_{\texttt{ref}}$) is tackled with an algorithm that return points satisfying only necessary optimality conditions.
Given a collection of partition sets of $\mathbb{Y}$, there is a flexibility in the way to index them. This choice of indexation impacts the shape of $\Phi$. Then, we may fine-tune the indexation of the partition sets of any partition of $\mathbb{Y}$ so that the resulting Reformulation ($\mathbf{P}_{\texttt{ref}}$) is solved as easily as possible.
Reformulation ($\mathbf{P}_{\texttt{ref}}$) is currently defined according to an extreme barrier [3], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible. We observed in [9, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value +$\infty$+ ∞. A progressive barrier [4] may be more efficient. We have initiated some research to associate an infeasibility metric to all points $x\in\mathbb{X}$ and use it in the progressive barrier. For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{ref}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). In all the problems we tested in [9, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is significantly easier with such a progressive barrier than than with an extreme barrier.
This framework may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for any $x\in X$). We plan to develop more advanced hybrid methods. Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly by optimizing jointly the partition set to consider and the point to consider into the set.
Finally, the class of composite problems illustrated in Section 4 is not the only class where the POf might be considered. A noticeable example lies in partially separable problems, where fixing some key variables makes the subproblem fully separable. Another example consists in problems where each variable has little individual influence on the objective value but their joint influence is much greater, similarly to the state variables in optimal control theory. As the theory is now stated, applying the POf requires only to follow the content of Section 3. We plan to reword our former work [10] accordingly.